# Is Mamba Capable of In-Context Learning?

### v1

@@ Abstract
This work provides empirical evidence that Mamba, a newly proposed selective structured state space model, has similar in-context learning (ICL) capabilities as transformers. We evaluated Mamba on tasks involving simple function approximation as well as more complex natural language processing problems. Our results demonstrate that across both categories of tasks, Mamba matches the performance of transformer models for ICL. Further analysis reveals that like transformers, Mamba appears to solve ICL problems by incrementally optimizing its internal representations. Overall, our work suggests that Mamba can be an efficient alternative to transformers for ICL tasks involving longer input sequences.
@@ 1 Introduction
Recent advancements in large-scale neural language modeling (Brown et al., 2020) have demonstrated that Transformer models (Vaswani et al., 2017) exhibit in-context learning (ICL) capabilities: after (self-supervised) pre-training, they can infer how to perform tasks only from input examples without the need for explicit training nor fine-tuning. This ability represents a departure from established in-weights learning of traditional machine learning and has sparked considerable academic interest. In contrast to meta-learning approaches (Hospedales et al., 2021), in transformer models ICL emerges from pre-training: without explicit training on a distribution of tasks and without having any specific inductive bias. Recent studies advanced the understanding of how transformers can implement and learn (variants of) in-context gradient descent when trained on distributions of simple supervised learning tasks, e.g., on linear regression tasks (Von Oswald et al., 2023; Ahn et al., 2023; Bai et al., 2023). Despite such results, whether pre-trained transformers perform in-context gradient descent on more complex tasks remains an ongoing discussion (Shen et al., 2023).
@@
Orthogonal to these investigations into the transformer architecture, recent work proposed deep state space models to overcome limitations of transformers in processing long sequences (Tay et al., 2021), such as S4 (Gu et al., 2021a) or H3 (Gu et al., 2021b). In essence, these models merge elements from recurrent and convolutional networks with state space approaches (Kalman, 1960). However, their success on NLP tasks was limited due to problems handling dense information tasks.
@@
This works conducts an investigation into the ICL capabilities of the recently proposed Mamba architecture (Gu & Dao, 2023), a successor to S4 and H3. Mamba has already shown its potential in different applications beyond NLP, such as (visual) representation learning (Zhu et al., 2024) or image segmentation (Ma et al., 2024). Concurrent to our work, the ICL capabilities of Mamba were investigated on synthetic language learning tasks by Akyürek et al. (2024). Our work is orthogonal to theirs, as we provide ICL results for both regression and NLP tasks, and additionally conduct a probing analysis on the representations at intermediate layers to better understand Mamba’s ICL solution mechanisms (for simple function classes).
Our analysis provides the following contributions to understanding Mamba’s ICL capabilities:
@@
  • We demonstrate that Mamba is capable of ICL and performs on-par with transformers in many cases This highlights Mamba as an efficient alternative to transformers for ICL tasks entailing longer sequences. In addition, we find that Mamba outperforms its predecessor S4, and a recent Recurrent Neural Network (RNN) architecture, RWKV (Peng et al., 2023).
@@
  • Using a simple probing approach, we provide preliminary insights into the mechanism by which Mamba incrementally solves ICL tasks. We find that the optimization processes exhibited by Mamba bear similarities to those of transformer models.
Outline  We begin by establishing Mamba’s ICL capabilities on simple function classes (Garg et al., 2022) in Section 2, comparing it to a causal transformer model in Section 2.1. Having established that Mamba is capable of ICL, we aim to understand its internal mechanisms to perform ICL in Section 2.2. To this end, we analyze how Mamba incrementally approximates the solution of the in-context task. Finally, we study Mamba’s ICL performance on more complex NLP in-context tasks from Hendel et al. (2023) in Section 3.
@@ 2 Investigation of Simple Function Classes
We followed the experimental protocol of Garg et al. (2022): each model is trained on a task distribution and then tested on the same distribution. This process was repeated for 3 regression task distributions, each falling into a specific function class: linear functions, 2-layer ReLU neural networks, and decision trees. Models trained on linear functions were also tested on out-of-distribution (OOD) tasks. Differently from Garg et al. (2022), we also tested the models on tasks with more input examples than seen during training, to test if they can extrapolate to longer inputs. Details on each task distribution are provided in Section A.2.3.
We compare Mamba to a causal transformer model using the GPT2 architecture (Radford et al., 2019), S4 and some baselines specific to each function class (see Section A.2.4). We intentionally compare to S4, a predecessor to Mamba, as it is a linear-time invariant model: its state dynamics are linear and constant through time. In contrast, Mamba is a linear time variant model: its state dynamics can change based on the current input. For architectural details on Mamba see Section A.1. To ensure a fair comparison, the architecture of Mamba and S4 is adjusted to have a comparable number of parameters to the transformer (9.5 M). For further training details we refer to Section A.2.1.
We removed the positional embedding used in the causal transformer by Garg et al. (2022), since we observed that it hinders the transformer’s ability to extrapolate to longer inputs, as also shown by Press et al. (2021). Similar to Müller et al. (2021), we argue this to be more natural in the present setup as the input is actually a set of data points, not a sequence.
To sample the inputs of each training task for linear regression, Garg et al. (2022) used a normal distribution, while we used a skewed Gaussian distribution. Our findings show that this improves robustness w.r.t. OOD tasks of both Mamba and transformers. Formally, we trained on linear functions of the form $F=\{f\,:\,f(x)=w^{T}x,w\in\mathbb{R}^{d}\}$ with input dimension $d=20$. For each training task, we sampled $w$ from an isotropic Gaussian $\mathcal{N}(0,I_{d})$ and $x_{1},\dots,x_{k},x_{k+1}$ from $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a skewed covariance matrix with its eigenbasis chosen at random and the $i$th eigenvalue proportional to $1/i^{2}$. Following this, we set $y_{i}=w^{T}x_{i}$ and assembled the input prompt $P=(x_{1},y_{1},x_{2},y_{2},...,x_{k},y_{k},x_{k+1})$. We used the mean squared error $\frac{1}{k+1}(\hat{y}_{i}-y_{i})^{2}$ as loss function, where $\hat{y}_{i}$ is the output of the model corresponding to $x_{i}$.
@@ 2.1 Analysis of in-distribution and out-of-distribution performance
(c) x-scale 0.333
(d) Random quadrants ————–
Figure 1: Comparative visualization of Mamba, S4, and transformer models trained on skewed linear regression and evaluated in-distribution (a) and out-of-distribution (b) - (d). The dotted vertical line indicates the number of in-context examples used for training.
Figure 1 shows our results for models trained on skewed linear regression tasks. Both Mamba and the transformer model are capable to solve the in-distribution and out-of-distributions tasks, while S4 is unable to do so. We find that Mamba’s ICL performance deteriorates more than that of the transformer as the number of in-context examples increases. We hypothesize that S4’s poor ICL performance is due to its linear time invariance. A similar hypothesis was also drawn by Gu & Dao (2023) for the task of selective copying. Besides that, both Mamba and Transformers exhibit input length extrapolation on noisy linear regression: the performance improves with more than 40 in-context examples.
Section A.2.5 provides further results for Mamba and the transformer model on decision trees and 2-layer ReLU neural networks regression tasks perform comparably when evaluated in-distribution (Figures 4(c) and 4(b)), while when trained on non-skewed linear regression tasks they are less robust to OOD tasks (Figure 6), even if they both perform well in-distribution (Figure 4(a)).
2.2 Mechanistic understanding via probing
(a) Skewed Linear Regression
@@ 2.2 Mechanistic understanding via probing
Figure 2: ICL ’Learning curves’ depicting the iterative optimization performance of Mamba and Transformer models on three regression tasks.
@@
To better understand how the Mamba and transformer models perform ICL, we aim to test if they both employ a solution strategy akin to iterative optimization, i.e., study if they incrementally improve their solutions layer after layer (Von Oswald et al., 2023; Ahn et al., 2023; Bai et al., 2023).
@@
We adopted a probing strategy similar to the one by (Geva et al., 2021) for transformer language models. Differently from other probing strategies like the one by Akyürek et al. (2023), who learn a (non-linear) probe on the high-dimensional intermediate representations, this strategy first passes the intermediate representations through the decoder at the end of transformer model used to map from the embedding to the output space (see Section A.2.6). The decoding procedure of Mamba is similar to the transformer, but includes a layer norm. To account for possible differences in scale and shift, we estimated a linear probing model on the in-context example data using least squares between the predictions and the regression target. We argue this to be a less biased probing strategy, since in our setup it reduces the degrees of freedom of the probe to just $2$ parameters per task: scale and shift.
@@
We conducted the probing analysis on transformers and Mamba models trained on skewed linear regression, ReLU neural networks, and decision trees. For skewed linear regression, we additionally compared to Gradient Descent (GD) and ${}^{++}$ as done in Von Oswald et al. (2023). ${}^{++}$ is a version of preconditioned gradient descent in which the data samples undergo a transformation ($x_{i}\leftarrow(I-\gamma XX^{T})x_{i}$) and we tuned $\gamma$ for optimal performance via a grid search (c.f., Figure 7 in Appendix), while the step-size was set for each task to the theoretically optimal value, i.e., as $2/(L+\mu)$, where $L,\mu$ are the largest and smallest eigenvalues of the empirical covariance matrix of the task.color=violet!20,]RG: Is this true?, JS: Yes We ran GD and ${}^{++}$ for 24 iterations to match the 24 layers of our Mamba model.
Figure 2 provides strong evidence to the hypothesis that both Mamba and the transformer employ an iterative solution scheme on the linear regression task, since the log-MSE decreases (almost) linearly. For ReLU neural networks and decision trees, the results are more ambiguous and further investigations are required, since the error decreases more abruptly, albeit sometimes decreasing gradually, and the final error is still relatively high. We also note that intermediate predictions for all three tasks achieve similar MSE. The comparison between GD, ${}^{++}$, Mamba, and the transformer in Figure 1(a) reveals that ${}^{++}$ achieves a performance level similar to Mamba, while GD shows slower convergence also due to the tasks having skewed covariance. Interestingly, the transformer model outperforms ${}^{++}$.
Finally, we compared the scalar scale and shift estimated by the linear probing model for Mamba and transformer in Figure 8 in the Appendix. Interestingly, we find that the linear probing model only makes small adjustments for linear regression and ReLU neural network, with scales close to one and shifts close to zero.

### v2

@@ Abstract
State of the art foundation models such as GPT-4 perform surprisingly well at in-context learning (ICL), a variant of meta-learning concerning the learned ability to solve tasks during a neural network forward pass, exploiting contextual information provided as input to the model. This useful ability emerges as a side product of the foundation model’s massive pretraining. While transformer models are currently the state of the art in ICL, this work provides empirical evidence that Mamba, a newly proposed state space model which scales better than transformers w.r.t. the input sequence length, has similar ICL capabilities. We evaluated Mamba on tasks involving simple function approximation as well as more complex natural language processing problems. Our results demonstrate that, across both categories of tasks, Mamba closely matches the performance of transformer models for ICL. Further analysis reveals that, like transformers, Mamba appears to solve ICL problems by incrementally optimizing its internal representations. Overall, our work suggests that Mamba can be an efficient alternative to transformers for ICL tasks involving long input sequences. This is an exciting finding in meta-learning and may enable generalizations of in-context learned AutoML algorithms (like TabPFN or Optformer) to long input sequences. The code to reproduce our experiments is available at github.com/automl/is_mamba_capable_of_icl.
@@ 1  Introduction
Recent advancements in large-scale neural language modeling (Brown et al.,, 2020) have demonstrated that Transformer models (Vaswani et al.,, 2017) exhibit in-context learning (ICL) capabilities: after (self-supervised) pre-training, they can infer how to perform tasks only from input examples without the need for explicit training nor fine-tuning. This ability represents a departure from established in-weights learning of traditional machine learning and has sparked considerable academic interest as a new type of meta-learning. In contrast to standard meta-learning approaches (Hospedales et al.,, 2021), ICL emerges in transformer models from pre-training: without explicit training on a distribution of tasks, without bi-level optimization, and without any specific inductive bias. Recent studies advanced the understanding of how transformers can implement and learn in-context gradient-based methods when trained on distributions of simple supervised learning tasks, e.g., on linear regression tasks (Von Oswald et al.,, 2023; Ahn et al.,, 2023; Bai et al.,, 2023). Despite such results, whether pre-trained transformers perform in-context gradient methods on more complex tasks remains an ongoing discussion (Shen et al.,, 2023).
@@
Orthogonal to these investigations into the transformer architecture, recent work proposed deep state space models to overcome limitations of transformers in processing long sequences (Tay et al.,, 2021), such as S4 (Gu et al., 2021a, ) or H3 (Gu et al., 2021b, ). These models merge elements from recurrent and convolutional networks with state space approaches (Kalman,, 1960). However, their success on NLP tasks was limited due to problems handling dense information tasks. A key feature of state space models is that they can run the forward pass in two modes with different complexity w.r.t. the input sequence length: a parallel mode, ideal for training but with superlinear time complexity and a recurrent mode with linear time complexity and more suited for inference. In contrast, the forward pass of transformer models has quadratic time complexity and hence it is less efficient.
@@
This work conducts an investigation into the ICL capabilities of the recently proposed Mamba architecture (Gu and Dao,, 2023), a successor to S4 and H3. Mamba has already shown its potential in different applications beyond NLP, such as visual representation learning (Zhu et al.,, 2024) or image segmentation (Ma et al.,, 2024). Concurrent to our work, the ICL capabilities of Mamba were investigated on synthetic language learning tasks by Akyürek et al., (2024) and on other tasks, including simple function classes, by Park et al., (2024). Differently from those works, we study the performance of the pre-trained Mamba model on NLP tasks, conduct a probing analysis on the representations at intermediate layers to better understand Mamba’s ICL solution mechanisms, and measuring the models’ ability to extrapolate beyond the context length used for training on simple function classes.
Models capable of ICL have shown good AutoML performance. A remarkable example are Prior-data Fitted Networks (PFNs) (Müller et al.,, 2022), which are (usually transformer) models pretrained on a range of synthetic tasks generated from a prior distribution, to approximate the posterior predictive distribution. Interestingly, by harnessing ICL and with a rich enough prior, PFNs, specifically TabPFN (Hollmann et al.,, 2023), exhibit strong generalization performance on real tabular data, enabling predictions on new data without the need for explicit specification and optimization of a table-specific model. Other applications of ICL relevant to AutoML are in Black-Box Optimization (Müller et al.,, 2023; Yang et al., 2023a, ; Chen et al.,, 2022), Learning-Curve Extrapolation (Adriaensen et al.,, 2024), and Time-Series Forecasting (Dooley et al.,, 2024). Our experimental setup for simple function classes, which closely follows the one in (Garg et al.,, 2022), is similar to that of PFNs, but the priors we use concern only simple regression tasks and the model only outputs the mean of the posterior predictive distribution.
Contributions
After introducing state-space models and Mamba more formally (Section 2), we make the following contributions:
@@
  • We demonstrate that Mamba is capable of ICL and performs on-par with transformers on simple function classes (Section 3.1) and more complex NLP tasks (Section 4). This highlights Mamba as an efficient alternative to transformers for ICL tasks entailing long sequences. In addition, we find that Mamba outperforms its predecessor S4, and RWKV (Peng et al.,, 2023), a recent parallel/recurrent architecture.
@@
  • Using a simple probing approach, we provide preliminary insights into the mechanism by which Mamba incrementally solves ICL tasks (Section 3.2). We find that the optimization processes exhibited by Mamba are similar to those of transformer models.
2  State Space Models and the Mamba Architecture
[…]
A fundamental limit of linear time invariant SSMs is that they do not have a mechanism to select which information to retain in the latent state based on the input sequence, and this makes them perform badly in tasks involving content-aware reasoning like selective copying, in which transformer models excel. The Mamba architecture (Gu and Dao,, 2023) is a linear time varying SSM which overcomes this limit via a selection mechanism allowing the latent state dynamics to change with the current input, thus enabling selective retention of information. In particular, state and outputs of Mamba follow the recursion (starting from $h_{0}=0$)
$h_{t}$ $=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t}$
$y_{t}$ $=C_{t}h_{t}$
where $\bar{A}_{t}=\exp(A\Delta_{t}),\bar{B}_{t}=\exp(A\Delta_{t})^{-1}\exp(\Delta_{t}A-I)\Delta_{t}B_{t}$ (zero order hold discretization), $B_{t}=W_{1}x_{t}+b_{1},C_{t}=W_{2}x_{t}+b_{2}$, $\Delta_{t}=\mathrm{softplus}(W_{3}x_{t}+b_{3})$ and $(A,W_{1},W_{2},W_{3},b_{1},b_{2},b_{3})$ are learnable parameters and $A$ is diagonal. The now time-varying discretization parameter $\Delta_{t}\in\mathbb{R}$ enables a gating mechanism which can selectively ignore the current input or reset the state. However, the selection mechanism hinders the use of convolutions for fast and parallel training. Despite this, Mamba has a similar training time as linear time invariant SSMs thanks to a hardware-aware algorithm that uses a parallel scan ($O(L)$ time complexity with only $O(\log(L))$ sequential steps) in place of FFT and stores and updates the large latent states only in the fast SRAM, rather than in GPU HBM memory. We refer to Gu and Dao, (2023) for further details on Mamba’s architecture.
@@ 3  Investigation of Simple Function Classes
In this section, we assess Mamba’s ability to learn task distributions involving simple function classes. We followed the experimental protocol of Garg et al., (2022): each model is trained on a task distribution and then tested on the same distribution. This process was repeated for 4 regression task distributions, each falling into a specific function class: linear functions, sparse linear functions, 2-layer ReLU neural networks, and Decision trees. Models trained on linear functions were also tested on out-of-distribution (OOD) tasks. Differently from Garg et al., (2022), we also tested the models on tasks with more input examples than seen during training, to measure whether they can extrapolate to longer inputs. Details on each task distribution are provided in Section A.1.3.
We compare Mamba to a causal transformer model using the GPT2 architecture (Radford et al.,, 2019) and to some baselines specific to each function class (see Section A.1.4). We also compare with S4, a linear time invariant model, to measure the impact of the selection mechanism of Mamba. To ensure a fair comparison, the architectures of Mamba and S4 are adjusted to have a comparable number of parameters to the transformer (9.5 M). For further training details, we refer to Section A.1.1.
We removed the positional encoding used in the causal transformer by Garg et al., (2022), since we observed that it hinders the transformer’s ability to extrapolate to longer inputs, as also shown by Press et al., (2021). Similar to Müller et al., (2022), we argue this to be more natural in the present setup as the input is actually a set of data points, not a sequence.
To sample the inputs of each training task for linear regression, Garg et al., (2022) used a normal distribution, while we used a skewed Gaussian distribution. Our findings show that this improves robustness w.r.t. OOD tasks of both Mamba and transformers. Formally, we trained on linear functions in the set $F=\{f\,:\,f(x)=w^{T}x,w\in\mathbb{R}\}$ with input dimension $d=20$. For each training task, we sampled $w$ from an isotropic Gaussian $\mathcal{N}(0,I_{d})$ and $x_{1},\dots,x_{k},x_{k+1}$ from $\mathcal{N}(0,\Sigma)$, where $\Sigma$ is a skewed covariance matrix with its eigenbasis chosen at random and the $i$th eigenvalue proportional to $1/i^{2}$. Following this, we set $y_{i}=w^{T}x_{i}$ and assembled the input prompt $P=(x_{1},y_{1},x_{2},y_{2},...,x_{k},y_{k},x_{k+1})$. We used the mean squared error (MSE) $\frac{1}{k+1}\sum_{i=1}^{k+1}(\hat{y}_{i}-y_{i})^{2}$ as loss function, where $\hat{y}_{i}$ is the output of the model corresponding to $x_{i}$. For all results in this section, we report the MSE/$d$ ($d=20$), which is what is reported by Garg et al., (2022) for linear regression.
@@ 3.1  Analysis of in-distribution and out-of-distribution performance
(c) Different orthants ————–
(d) x-scale 0.333
(e) y-scale 0.333
(f) $d/2$-dimensional subspace ————–
Figure 1: Comparative visualization of Mamba, S4, and transformer models (3 training seeds per method) trained on skewed linear regression and evaluated in-distribution (a) and out-of-distribution (b) - (f). The dashed vertical line indicates the number of in-context examples used for training (40).
(a) Sparse LR
@@
Figure 2: Comparison of Mamba and transformer models (3 training seeds per method) trained and tested on the same task distribution. The dotted vertical line indicates the number of in-context examples used for training (40 for sparse linear regression, 100 for ReLU neural network and Decision tree).
Results are shown in Figures 1 and 2. In skewed linear regression, both Mamba and the Transformer model closely match the least squares baseline both in-distribution and out-of-distribution as long as the context length is shorter than the one used for training. Increasing the context length generally decreases performance, by an amount depending on the training run: two runs of the transfomer model present very little degradation while Mamba degrades substantially for all three runs. In contrast, S4 performs much worse than the least squares baseline in all setups. We hypothesize that S4’s poor ICL performance is due to its linear time invariance. A similar hypothesis was also drawn by Gu and Dao, (2023) for the task of selective copying. Similar results for Mamba and the Transformer also hold for sparse linear regression, while for neural networks and Decision trees, Mamba and the transformers are comparable and even show a promising input length extrapolation: the performance improves after the dotted vertical line indicating the number of in-context examples used during training. Interestingly, in the case of Decision trees, two out of three runs of the Transformer perform substantially better than Mamba, while Mamba’s error degrades less than the one of the Transformer for ReLU neural networks. We also note that when trained on non-skewed linear regression tasks, Mamba and Transformers are less robust to OOD tasks (Figure 9), even if they both perform well in-distribution (Figure 8). In the appendix we also report the performance when increasing the number of in-context examples during training for skewed linear regression (Figure 10).
3.2  Mechanistic understanding via probing
(a) ReLU NN
(b) Decision tree
(c) Sparse LR
(d) Skewed LR ——————-
Figure 3: ICL ’Learning curves’ depicting the iterative optimization performance of Mamba and Transformer models on three regression tasks. Ratio of layers is the layer index divided by the total number of layers for each model (12 for the Transformer and 24 for Mamba). ReLU NN and Decision tree evaluated at $k=100$ in-context examples and sparse/skewed LR at $k=40$ in-context examples (i.e. the same number of in-context examples used during training).
(a) Mamba
(b) Transformer ——————-
@@ 3.2  Mechanistic understanding via probing
Figure 4: ICL ’Learning curves’ depicting the iterative optimization performance of Mamba and Transformer models (both trained with 40 in-context examples) on skewed linear regression when tested on varying the number of in-context examples from 40 to 160.
@@
To better understand how the Mamba and transformer models perform ICL, we aim to test if they both employ a solution strategy akin to iterative optimization, i.e., we study if they incrementally improve their solutions layer after layer (Von Oswald et al.,, 2023; Ahn et al.,, 2023; Bai et al.,, 2023).
@@
We adopted a probing strategy similar to the one by Geva et al., (2021) for transformer language models. Differently from other probing strategies like the one by Akyürek et al., (2023), who learn a (non-linear) probe on the high-dimensional intermediate representations, this strategy learns a linear probe on the output of the decoder applied to the intermediate representations after each layer. We argue this to be a less biased probing strategy, since in our setup it reduces the degrees of freedom of the probe to just $2$ parameters per task (scale and shift), since the models have scalar output.
More specifically, our probing strategy works as follows. Let $\{(x_{i},y_{i})\}_{i}$ be i.i.d. samples belonging to one task. To compute and evaluate intermediate predictions, we use $k$, $m$ and $n$ examples for the train, validation, and test set, respectively. First, we separately feed the prompts $(x_{1},y_{1},x_{2},y_{2},...,x_{k},y_{k},x_{j})$ for $j=k+1,\dots,k+m+n$ to the model, to obtain for each token $x_{i}$ the internal representations $z_{l,i}$ for each layer $l$ and the final prediction $\hat{y}_{i}=g(z_{L,i})$, where $g$ is the decoder and $L$ is the index of the last layer. Then for each layer $l$ and each token $i>k$, we first obtain intermediate scalar predictions $\tilde{y}_{l,i}=g(z_{l,i})$. Since $g$ is not meant to be used on intermediate representations, we finally adjust the predictions by computing $\hat{y}_{l,i}=a_{l}g(z_{l,i})+b_{l}$, where the scalar scale and shift parameters $a_{l}$ and $b_{l}$ are obtained by least squares on $(\tilde{y}_{l,k+1},y_{k+1})\dots(\tilde{y}_{l,k+m},y_{k+m})$ (i.e., using the validation tokens). We can then measure the accuracy of these intermediate predictions with the normalized mean squared error on the test tokens $(nd)^{-1}\sum_{j=1}^{n}(\hat{y}_{l,k+m+j}-y_{k+m+j})^{2}$ ($d=20$). We ran our analysis over 128 tasks (e.g. different linear regression weight vectors) with $k$ as specified in Figures 3 and 4, $m$=10 and $n$=40. We conducted the probing analysis on Mamba and transformer models trained on skewed linear regression, sparse linear regression, ReLU neural networks, and Decision trees.
Figure 5: Skewed LR
@@
For skewed linear regression, we additionally compared to Gradient Descent (GD) and GD++ as done in Von Oswald et al., (2023). GD++ is a version of preconditioned gradient descent in which the data samples undergo a transformation ($x_{i}\leftarrow(I-\gamma XX^{T})x_{i}$) and we tuned $\gamma$ for optimal performance via a grid search (c.f., Figure 11 in Appendix), while the step-size was set for each task to the theoretically optimal value, i.e., as $2/(L+\mu)$, where $L,\mu$ are the largest and smallest eigenvalues of the empirical covariance matrix of the task. We ran GD and GD++ for 24 iterations to match the 24 layers of our Mamba model. Figure 3 provides strong evidence to the hypothesis that both Mamba and the transformer employ an iterative solution scheme on the skewed and sparse linear regression task (Figure 3(d)), since the log-MSE decreases (almost) linearly. In ReLU neural networks, the error also decreases somewhat gradually after some layers. Interestingly, for Decision trees the error stays high for both models for more than half of the initial layers. The comparison between GD, GD++, Mamba, and the transformer in Figure 3(d) reveals that both GD and GD++ are outperformed by Mamba and the Transformer, while GD converges more slowly due to the tasks having skewed covariance.
In Figure 4 we show how, similarly to Figure 1(a), having more in-context examples than the ones used during training negatively affects the learning curves for linear regression. Both models never seem to take advantage of the additional number of examples in this task: the performance with 40 in context examples (the number used for training) is the best (or close) for all layers. Moreover, we note that the learning curve for Mamba with 160 in-context examples exhibits a clear U-shape: the model actually found a good solution in intermediate layers but ultimately degrades.
In Figure 5, we compare the scale and shift parameters estimated by the linear probing model for Mamba and the transformer in the context of Skewed Linear Regression. Additionally, for sparse linear regression, ReLU neural networks, and Decision trees, we present the findings in Figure 12 in the Appendix. Interestingly, for linear regression and ReLU neural networks, the linear probing model applies only minor modifications, with scale values near one and shift values approaching zero after the first layers. We also observe that for all tasks, the variance across tasks of the estimated scale and shift is higher in the early layers of the model and quickly approaches zero as we move towards the output layer.